OpenStudy (anonymous):

what numbers when multiplied equals -30 but when added together equals -4

OpenStudy (anonymous):

factors of -30 are (1 and -30), (-1 and 30), (5 and -6), (-5 and 6), (3 and -10), (-3 and 10), (2 and -15), (-2 and 15) none of which when added, equal -4.

OpenStudy (anonymous):

so then how am i supposed to finish this question

OpenStudy (anonymous):

there are no solutions

OpenStudy (anonymous):

you say there is (no solution)

OpenStudy (anonymous):

there has to be because it wont accept no solution or zero as an answer the original problem is Write the function in factored form check by multiplication y=-4x^3-4x^2+120x

OpenStudy (anonymous):

the final answer its saying should look like y=-4x(x+?)(x-?)

OpenStudy (anonymous):

what was your work to factor the problem

OpenStudy (anonymous):

Idk i jusr crumpled up the paper and threw it math is pissing me off today

OpenStudy (anonymous):

taking out a -4x, we get y=-4x(x^2+x-30). if you are setting y=0, then either x=0 or (-=x^2+x-30)=0. In which case we need something that when multiplied gets -30 and added gets 1, which would be -5 and 6. y=-4x(x-5)(x+6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks